OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (geometry_hater):

you need to get y by itself first

OpenStudy (anonymous):

hi michael, please try to arrange it y=mx+b

OpenStudy (geometry_hater):

So id subtract 2x tot he other side then divide by -5

OpenStudy (anonymous):

I think it's going to be like 2/5.. does that make since

OpenStudy (anonymous):

\[y = {2x + 10 \over 5}\] then, \[y = (\frac{ 2 }{ 5 }x) + 2\]

OpenStudy (anonymous):

m = 2/5 b = 2

OpenStudy (anonymous):

the answer is 2? not 2/5

OpenStudy (anonymous):

if you are getting for the slope "m" it should be (2/5) yes, but if the question is find the y-intercept "b" then that's the time that you should be getting an answer of 2. I hope this helps.

OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10? negative five halves five halves two fifths negative two fifths none of the above

OpenStudy (anonymous):

two fifths